OpenStudy (anonymous):

I know how to solve simultaneous equation but i think for this question i dont know some rule. X+2y=2 x-y=-2 Eliminate method 1st eq -2nd eq 3y=4 y=4/3 What i dont get is how come in solution they dont multiply 2nd eq by 2 when i do it i get right x value but wrong y value

OpenStudy (haseeb96):

just first multiply 2nd equation by 2 and then use elimination method

OpenStudy (haseeb96):

and first question also by 2 2x+4y=4 2x-2y=-4 ----------- +6y =8 y=8/6 y=4/3

OpenStudy (anonymous):

I know but in solution they didnt do it what my question is If the question is like x-y=9 X+2y=8 Is solve different to 2x-y=7 3x+y=6

OpenStudy (haseeb96):

now just put y=4/3 into any one equation either in 1 or 2 and get the x i put it in equation 1 2x+4y=4 2x+4(4/3) =4 6x +16 /3 =4 6x +16 =4 x 3 6x +16 =12 6x=12-16 6x=-4 x=-4/6 x=-2/3

OpenStudy (haseeb96):

2x-y =7 eq#1 3x+y =6 eq#2 since just use the addition method so ,it will be ------- 5x =13 x=13/5 and now put x into any equation and get the y

OpenStudy (anonymous):

U know the initial step 2x+4=4 +2x-2y=-4 How come 4+(-4) became 8 do we neglect the sign

OpenStudy (anonymous):

Srrry after long time i have started doing maths

OpenStudy (haseeb96):

in the elimination method sign of second terms change if u are using elimination method if there is + ve sign then it will convert into -ve okay?

OpenStudy (haseeb96):



OpenStudy (haseeb96):

have u got it?

OpenStudy (anonymous):

Thank u

OpenStudy (haseeb96):

not say thank u Just help three persons and say to them help more three persons anywhere okay?

OpenStudy (anonymous):

@Haseeb96 taking movie morals too seriously aye !

OpenStudy (haseeb96):

@ayeshaafzal221 just be +ve I watch movies for both the things i.e. entertainment and also lesson it is good lesson for us if we adopt it but we are pakistani and t how can we keep our mouth silence or irritate the other till our words didnt hurt the other one 's heart and mind. :P

OpenStudy (anonymous):

I dont know what pakistan u talking about i aint like that k maybe a little bit but ur right we shuld adopt good things so i shall help three people!:D @Haseeb96

OpenStudy (haseeb96):

haha oh ho acting karna koi tm se sikhe :D

OpenStudy (haseeb96):

do u not know the pakistan ? jis mulk ne tmhe rehne ki ijazat di aj os ka name lete hwe tmhe shamram ati hai aur dosro ko bhata ne mein bhi

OpenStudy (anonymous):

K i dont know that much of urdu @Haseeb96

OpenStudy (haseeb96):

okay where are u from?